40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2464     Entry time: Tue Dec 29 04:28:27 2009
Author: kiwamu, rana, haixing 
Type: Update 
Category: Cameras 
Subject: New Video Switch Installed 

We have installed the new Video Matrix.

Its still in an intermediate state, so don't try to "fix" anything before Kiwamu and I get back onto it in the afternoon.

The status so far is that we have removed the old switch (it was a 256 input x 128 output !! mux)  and installed the new one in the same rack. We have hooked it up to the CDS network and have set up its matrix by using the web interface (i.e. NOT EPICS).


Along the way, we discovered that there is lack of impedance matching in the video all over the 40m. Video signals are RF and need to be treated that way. The PSL signals are T'd around and sent on 50 Ohm cables to high impedance monitor inputs.

We should eliminate any switches besides the new one (called Luciana) and control the PSL's Video Monitor from the main MUX interface. No more Rogue Video Switches.

 


Another couple of things we have found is about RCR camera.

(1) The long cable which connects the RCR camera box and the video matrix doesn't work. Although the signal is alive and we can see it by the local tv monitor nearby PSL.

(1) The reflected beam going to the camera is too weak to see in the monitor.  We found a strange polarized cube splitter in front of the camera. We should modify it sooner or later.

ELOG V3.1.3-